OpenStudy (anonymous):

this is probably a basic question but i need help with the formula to find the longest length of the triangle? One side is 8 on side is 11 and the other is x

OpenStudy (anonymous):

a^2 + b^2 = c^2

OpenStudy (anonymous):

8^2 = 64 11^2 = 121 121+ 64 = 185

OpenStudy (anonymous):

so it is just as simple as adding the side together and squaring them?

OpenStudy (anonymous):

woopsies 11 would be the long side so we are solving for b.

OpenStudy (anonymous):

the long side is the one im solving for

OpenStudy (anonymous):

Pretty much. you have to use the Pythagorean theorem

OpenStudy (anonymous):

so then would i square 185 to get the answer

OpenStudy (anonymous):

Then the long side is c which is your hypotenuse and a and be are your legs

OpenStudy (anonymous):

185^2?

OpenStudy (anonymous):

I think you have to do 360 - 185 which gives you 175, then you √175 and it gives you 13.22. I'm not sure if that is the answer though.

OpenStudy (anonymous):

so round to the nearest whole number which would be 13, and see